The length of a rectangle is 4cm more than its width. Its diagonal is 8cm more than its width. What is the width of the rectangle? Show the solution.
Answer by mananth(16949)

.
(x+8)^2 = x^2 +(x+4)^2 ( Pythagoras theorem)
x^2 +16x +64 = x^2 +x^2 +8x +16
Simplify
x^2-8x -48 =0
(x-12)(x+4)=0
x= 12 or -4
Neglect negative
width =12 cm